urgent can you please help me out by telling some substitutions that can be done to solve indefinite integration

urgent


can you please help me out by telling some substitutions that can be done to solve indefinite integration


 

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

For terms of the form x2 + a2 or √x2 + a2, put x = a tanθ or a cotθ

 

          For terms of the form x2 - a2 or √x2 – a2 , put x = a sec θ or a cosecθ

 

          For terms of the form a2 - x2 or √x2 + a2, put x = a sin θ or a cosθ

 

•         If both √a+x, √a–x,  are present, then put x = a cos θ.

 

•         For the type √(x–a)(b–x), put x = a cos2θ + b sin2θ

Win exciting gifts by answering the questions on Discussion Forum.

So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Sagar Singh